OpenStudy (anonymous):

Graph the functions. 1) y=|x+2|-1 2) y=-4|x|

OpenStudy (anonymous):

Do you know how to graph y = |x|?

OpenStudy (anonymous):

I don't

OpenStudy (anonymous):

Ok you can get the graphs you want from the simpler graph y = |x|

OpenStudy (anonymous):

Maybe If I see it, haven't graphed in a while..

OpenStudy (anonymous):

the graph of y= |x| looks like a V. with the vertex at the origin.

OpenStudy (accessdenied):



OpenStudy (accessdenied):

sorry that's y=|x|

OpenStudy (anonymous):

ah.

OpenStudy (anonymous):

How would I graph first equation then?

OpenStudy (accessdenied):

first equation is a transformation of the parent function y= |x|, or more specifically shifting the graph / changing its position. General shifts of |x|: y = |x - h| + k h units to the right, k units up In this case, h is -2 (x+2 = x-(-2)) and k = -1 So, the graph would look something like...